OpenStudy (anonymous):

how do you add 5 radical 6 +radical 6

jimthompson5910 (jim_thompson5910):

Think of 'radical 6' as a variable let's call that variable 'z' for now

jimthompson5910 (jim_thompson5910):

so 5 radical 6 +radical 6 is the same as 5z + z

jimthompson5910 (jim_thompson5910):

5z + z = ??

OpenStudy (anonymous):

5z squared?

jimthompson5910 (jim_thompson5910):

5z + z = 5z + 1z = (5+1)z = (6)z = 6z

jimthompson5910 (jim_thompson5910):

so 5z + z = 6z

OpenStudy (anonymous):

6z

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now replace z with what we made it equal to..which was radical 6

jimthompson5910 (jim_thompson5910):

so the answer is 6 radical 6 or \[\large 6\sqrt{6}\]

OpenStudy (anonymous):

okay thanks (:

jimthompson5910 (jim_thompson5910):

you're welcome